please i need help on these question can any one help me????
please help......
1)Find the 111th term of an arithmetic sequence that has a first term of an arithmetic sequence that has a first term of 4 and a common difference of -0.8
2)Write the following series in summation notation. Use i as index variable and start I at 1 2+4+6+8+…+124
3)Find the number of terms of a geometric sequence with, first term1/64, common ratio 2, and last term 512
4)What is the coefficient of a^3 z^3 in the expansion of (a+z)^12

Answer by anantha(86)

solution for 1st question
sol:
1st term in arithematic sequence (a)=4
commom difference (d)= -0.8
the nth term in arithematic sequence is
tn= a+(n-1)*d
now we find out the 11th term in arithematic sequence
n=11
a=4
d= -0.8
substitute these values in the above formula we get
t11= 4+(11-1)*(-0.8)
t11=4+(10)*(-0.8)
t11=4-8
t11= -4
11 th term in arithematic sequence is -4